OpenStudy (cutepuppies):

Which statements about the value of the digits in the number 8,888,888 are true? Choose all answers that are correct. A. The digit in the hundred thousands place is 1/10 the value of the digit in the millions place. B. The digit in the thousands place is 1/8 the value of the digit in the ten thousands. C. The digit in the ten thousands place is 10 times the value of the digit in the thousands place. D. The digit in the hundreds place is 80 times the value of the digit in the tens place.

OpenStudy (cutepuppies):

@Willie579

OpenStudy (cutepuppies):

@girlmeetsworld

OpenStudy (girlmeetsworld):

I don't know this. ;-; Sorry.

OpenStudy (cutepuppies):

oh its ok

OpenStudy (girlmeetsworld):

Good luck.

OpenStudy (cutepuppies):

thank you

OpenStudy (cutepuppies):

can anyone help me plz

OpenStudy (anonymous):

first one is true, since 1 000 000 /10 = 100 000, that is, if the digit is the same

OpenStudy (cutepuppies):

so its a

OpenStudy (anonymous):

but not only a, it says all answers, so there is more than one what do you think, going by this logic, is also correct out of those three left answers?

OpenStudy (cutepuppies):

ok do u know the rest of the answers

OpenStudy (anonymous):

yes, but if i simply tell you you wont learn anything, so i wouldnt be really helping you

OpenStudy (cutepuppies):

my big sister knows this but she will explian later so its ok you can give me direct answers

OpenStudy (anonymous):

ok, this is system with base 10, so the answer is c... if the base was 8, other answers might be possible, but in this case, it's a and c

OpenStudy (cutepuppies):

ok thank you very much for your help

OpenStudy (anonymous):

here is how it's done, for a 8 000 000 /10 = 800 000 and for c, 8000*10 = 80000

OpenStudy (anonymous):

glad i've been able to be helpful

OpenStudy (cutepuppies):

ok got it